OpenStudy (anonymous):

caroline runs 3 1/2 miles in 22 1/4 minutes. how many miles does she run per minute? A. 3/22 mile per minute B. 62/221mile per minute C. 14/89 mile per minute D. 4/39 mile per minute

OpenStudy (jack1):

caroline runs 3 1/2 miles in 22 1/4 minutes speed = \(\huge \frac {3.5~~ miles}{22.25~~minutes}\) so 3.5 miles in 22.25 minutes... how far in 1 minute?

OpenStudy (anonymous):

B

OpenStudy (jack1):

no

OpenStudy (jack1):

why b?

OpenStudy (jack1):

multiply top and bottom by 4, so u get whole numbers

OpenStudy (anonymous):

OH I think it D because I divided all of them and D = 9.75

OpenStudy (jack1):

ok... how does 9.75 come into this tho?

OpenStudy (anonymous):

I don't really know

OpenStudy (jack1):

ok... so get \(\huge \frac {3.5~~ miles}{22.25~~minutes}\) and times top and bottom by 4... what do u get?

OpenStudy (anonymous):

3.5 = 14 22.25=89

OpenStudy (jack1):

perfect

OpenStudy (jack1):

so answer is...?

OpenStudy (anonymous):

C do you mind if I fan you

OpenStudy (jack1):

yep, C and fan away dude, i never mind that hey ;)